If the length of a simple pendulum is doubled keeping its amplitude constant its energy will be

a) 

Unchanged

b) 

Doubled

c) 

Four times

d) 

Halved

 

Grade:12

3 Answers

Arun
4 years ago

 

As we learnt in

Time period of oscillation of simple pendulum -

T=2\pi \sqrt{\frac{l}{g}}

- wherein

l = length of pendulum 

g = acceleration due to gravity.

 

 Halved

 

Hence option D is correct
Sanjeev Narang
4 years ago
Simple pendulum is doing simple harmonic motion . so, Total energy is given by:
      T. E. = 1/2*m*A2*(\omega)2 
But (\omega)2 =g/l
Then, T. E. =1/2*m*A2*g/l
As mass(m), 1/2, Amplitude(A), and Acceleration due to gravity(g) are constant, then:
         T. E. is inversely proportional to length of the simple pendulum(l). Therefore if length become doubled T.E. becomes halved.
So option (d) is the answer. 
 
 
Khimraj
4 years ago
 
the answer is (d) halved. we know the formula t = 2pi * root(l/g).
so when l’ = 2l. t becomes root(2) times more.